A 1500 kg car is moving westward with a acceleration of 2 m/s2 and a 3000 kg track is travelling east with a acceleration of 1.6 m/s2 from the same position starting from the rest. Find the position, velocity and acceleration of the center of mass of the system from their starting point after 1 hour.
1
Expert's answer
2012-03-16T11:29:28-0400
Xcar=Vcar*t=(Vcar0+acart)t=acar*t^2=2*(3600*3600)=25920000m=25920km
Xtrack=1.6*3600*3600=20736000m=20736km
Xcentr=(Xcar*mcar+Xtrack*mtrac)/(mcar+mtrac)=22464km
Vcentr=(Vcar*mcar+Vtrack*mtrac)/(mcar+mtrac)=(acar*t*mcar+atrack*t*mtrac)/(mcar+mtrac)=6240m/s
acetnr=(acar*mcar+atrack*mtrac)/(mcar+mtrac)=52/3m/s^2